OpenStudy (anonymous):

If a5/a3=4/9 and a2=4/9 find the nth geometric sequence.?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ \dfrac{a_5}{a_3}=\dfrac{4}{9} \\~\\ \dfrac{ar^4}{ar^2}=\dfrac{4}{9} \\~\\ r^2=\dfrac{4}{9} \\~\\ r=\sqrt{\dfrac{4}{9}} \\~\\ r=\dfrac{2}{3}\\~\\~\\ a_2=\dfrac{4}{9}\\~\\ ar=\dfrac{4}{9}\\~\\ a\times \dfrac{2}{3}=\dfrac{4}{9}\\~\\ a=\dfrac{4}{9}\times \dfrac{3}{2}\\~\\ a=\dfrac{2}{3}\\~\\~\\ \text{nth term}\\ =ar^{n-1}\\~\\ =\dfrac{2}{3}\left(\dfrac{2}{3}\right )^{n-1}\\~\\ }\end{align}\)